If the total hanging massif 50g increased 4 times without changing the spring constant, how would this affect period of oscillation and/or the frequency?
1
Expert's answer
2019-10-30T12:31:44-0400

The formula for the period of oscillation:


"T=2\pi\sqrt{\frac{m}{k}}"

Thus, for the period:

"\frac{T'}{T}=\frac{2\pi\sqrt{\frac{4m}{k}}}{2\pi\sqrt{\frac{m}{k}}}=2"

Therefore,


"\frac{f'}{f}=\frac{T}{T'}=\frac{1}{2}"